OpenStudy (anonymous):

y(4)-4y"=t^2 +e^t, please help

OpenStudy (anonymous):

I got stuck when calculating the partial solution, \[Y_{p1}= At^2 +Bt+C\]

OpenStudy (anonymous):

the homogeneous solution is \[y_g= C_1+C_2t+C_3e^{2t}+C_4e^{-2t}\]It's ok,

OpenStudy (anonymous):

At^2+bt+C+De^t should be your solution

OpenStudy (anonymous):

I separate the partial solution into 2 part, you are right with De^t I just calculate the polynomial part and get stuck when second derivative is a constant.

OpenStudy (anonymous):

\[Y_{p1}=At^2 +Bt +C\\Y'_{p1}=2At+B\\Y"_{p1}=2A\\Y"'_{p1}=0\] then \[-4(2A) = t^2\]it doesn't make sense at all

OpenStudy (anonymous):

@bahroom

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

@loser66 there is try solving the equation with \(z=y''\)

OpenStudy (anonymous):

if so, I have y_p= -1/4t^2-1/8

OpenStudy (anonymous):

how to plug back?

OpenStudy (anonymous):

$$z''-4z=0$$Let \(z=e^{\lambda t}\):$$\lambda^2e^{\lambda t}-4e^{\lambda t}=0\\e^{\lambda t}(\lambda^2-4)=0\\\lambda =\pm2$$hence \(z_c=C_1e^{2t}+C_2e^{-2t}\) is the general solution for our homogeneous equation. For our particular solution, observe we need to come up with \(t^2,e^t\) terms on our right-hand side so assume a solution of the form \(z_p=A+Bt+Ct^2+De^t\), yielding \(z_p''=2C+De^t\):$$2C+De^t-4(A+Bt+Ct^2+De^t)=t^2+e^t\\2C+De^t=4A+4Bt+(4C+1)t^2+(4D+1)e^t$$so clearly \(C=2A,D=4D+1,B=0,4C+1=0\) hence \(A=-1/8,B=0, C=-1/4,D=-1/3\) and \(z_p=-\frac18-\frac14t^2-\frac13e^t\). Observe, then, that we have the general solution \(z=z_c+z_p=C_1e^{2t}+C_2e^{-2t}-\frac13e^t-\frac14t^2-\frac18\). To determine the general solution for \(y\), observe:$$z=y''\\y=\int\int C_1e^{2t}+C_2e^{-2t}-\frac13e^t-\frac14t^2-\frac18\,dx\,dx=\dots$$

OpenStudy (anonymous):

The reason your method failed is because you guessed a nonsense particular solution. Observe that since we only have \(y'',y^{(4)}\) in our equation, they must contain \(t^2,e^t\) terms, yet your given particular solution "guess" gives \(y''=2A\) and \(y''=0\). Instead, you should've tried with \(y_p=A+Bt+Ct^2+Dt^3+Et^4+Fe^t\)... :-) that seemed ugly, though, so I rewrote it as a 'nicer' linear differential equation of lower order in \(z=y''\) so that you can deal with the ugliness later during integration.

OpenStudy (anonymous):

how can I know this method? why z = e^lambda t?

OpenStudy (anonymous):

This method is just basic substitution and the basic approach to solving inhomogeneous linear differential equations. Assuming \(z=e^{\lambda t}\) lets us find the fundamental solutions of our homogeneous complementary equation \(z''-z=0\). We then used the method of undetermined solutions to find a particular solution.

OpenStudy (anonymous):

without it, we still have lambda = +/- 2?

OpenStudy (anonymous):

I don't understand your question

OpenStudy (anonymous):

let see the equation after replace z = y" it's z"-4z = t^2 +e^t homogeneous part is z" -4z=0 characteristic equation is \(\lambda^2-4=0\\\lambda = \pm2\) which yield \[Z_g= C_1e^{-2t}+C_2e^{2t}\]

OpenStudy (anonymous):

Thanks a lot, You are only one person who can give out a neat solution like this. I appreciate your help

OpenStudy (anonymous):

@looser66 yes very good :-) that is indeed your general solution for the homogeneous equation \(z''-4z=0\).

OpenStudy (anonymous):

You get the exact same solution for \(y^{(4)}-4y''=0\), in fact :-) Let \(y=e^{\lambda t}\)$$\lambda^4e^{\lambda t}-4\lambda^2e^{\lambda t}=0\\\lambda^2e^{\lambda t}(\lambda^2-4)=0\\\lambda=0\,(\text{mult}\,2),\pm2$$Observe we have a repeated root hence \(y_c=C_1+C_2t+C_3e^{-2t}+C_4e^{2t}\). The solution for \(z\) is in face equivalent, i.e.:$$z=y''\\y=\int\int z\,dx\,dx=\int\int C_1e^{-2t}+C_2e^{2t}\,dx\,dx=\int-\frac12C_1e^{-2t}+\frac12C_2e^{2t}+C_3\,dx\\y=\frac14C_1e^{-2t}+\frac14C_2e^{2t}+C_3t+C_4$$... with an appropriate change in the labeling of constants :p

OpenStudy (anonymous):

got you. so, we must do that step? I mean let y = e^lambda t? and there are some more steps as you did ( take double integral ) to get the final answer?

OpenStudy (anonymous):

Ha !! It takes a long time to solve.

OpenStudy (anonymous):

well, with the \(z=y''\) substitution yes you must integrate to get back a solution \(y\) :-p but I think it is easier to substitute first than to do it the long way :p

OpenStudy (anonymous):

But anyways we do the \(e^{\lambda t}\) steps because it yields our characteristic equation! :p

OpenStudy (anonymous):

You could just memorize the form of the characteristic equation but where's the fun in that? :p more fun to derive

OpenStudy (anonymous):

NOOOOOO , it's not fun aaatttt aaaalll.

OpenStudy (anonymous):

I saw other person has the interesting problem, I tagged you there, come to help him, I want to see, too. PPleease

OpenStudy (anonymous):

differential equation, too.